OpenStudy (anonymous):

Let m denote the slope of the tangent to the curve y=X^3 - 4x^2 +x +4 at the point (-1 , -2) . If the slope of the tangent to the curve y=3x^2 + ax + b at (2 , 0) on the curve is m , dertermine the values of a and b.

OpenStudy (anonymous):

@Arnab09 @FoolForMath @experimentX @eliassaab @yash2651995

OpenStudy (lgbasallote):

are you allowed to use derivatives?

OpenStudy (anonymous):

yes

OpenStudy (yash2651995):

m=12? a=0? b=any value?

OpenStudy (yash2651995):

b will have a fixed value.. (2,0) should satisfy that equation

OpenStudy (yash2651995):

b=-12?

OpenStudy (anonymous):

@yash2651995 u r correct a=0 b=-12 how?????

OpenStudy (anonymous):

@yash2651995 can u show me wat u have done...

OpenStudy (yash2651995):

first equation.. f(x)=y=X^3 - 4x^2 +x +4 f'(x)=3x^2-8x+1 at (-1,-2) x=-1 so put its value here^ f'(x)=m i.e. instantaneous slope at that point that gives m=12 now differentiate second eq g(x)=y=3x^2 + ax + b g'(x)=6x+a =m at (x,y)= (2,0) (,given) 12+a=m=12 ie a=0 now g(x) has to satisfy (2,0) as the point will be on it as we are asked about slope on it.. now, y=3x^2 + ax + b a=0 x=2 y=0 means b=-12

OpenStudy (anonymous):

May i have help